# Probing arbitrary polarized photon pairs undergoing double Compton scatterings by a dedicated MC simulator validated with experimental data

### v1

@@ 5.4 Selection of high visibility region and detection efficiency corrections
Figure 10: The scattering angle distributions $\tilde{\Theta}_{a}~{}\text{vs}~{}\tilde{\Theta}_{b}$ of primary photons for (a) a perfect detector with 100% geometrical and detection efficiency, (b) the experimental data and (c) MC simulations of |$\ket{\psi^{+}}=\frac{1}{\sqrt{2}}(\ket{\text{HV}}+\ket{\text{VH}})$ state. The red circles represent the high-visibility area.
As aforementioned, the shape of the $\Delta\hat{\Phi}$ distribution for the |$\ket{\psi^{+}(\pm 45^{\circ})}$ initial state should be flat, and any systematic deviation from the flat line must be the manifestation of the detector-induced effects. Therefore, we use this MC sample to calculate the bin-per-bin correction coefficients $\alpha_{i}$ that would take into account the detector-induced deformation in the $\Delta\hat{\Phi}$ degree of freedom:
@@ 5.5 Correlation plots
The shape of the $\Delta\hat{\Phi}$ spectra can be described by $f(\Delta\hat{\Phi})=a(1+b\,\cos(2\Delta\hat{\Phi}))$ function, where $a$ represents the size of the quantum–uncorrelated part, while the $\mathcal{V}^{2}=|b|$ can be interpreted as a squared visibility factor introduced in  Eq. 7. The experimental data and MC simulations distributions are presented in Fig. 11. Both distributions are corrected for the detector-induced effects. The shapes of both data and MC distributions are in very good agreement. The weighted maximum likelihood technique was used to fit the $f(\Delta\hat{\Phi})$ distributions. The extracted fit parameters for MC and experimental data are equal within the obtained uncertainties. The determined value of squared visibility is equal to $\mathcal{V}^{2}_{exp}=0.27\pm 0.02$ and $\mathcal{V}^{2}_{MC}=0.28\pm 0.03$, for the experimental data and MC simulations, respectively. The quoted uncertainties are calculated based on the fitted parameter uncertainties and include only the statistical effects.
@@
Figure 11: The comparison of experimental and MC $\Delta\hat{\Phi}$ distributions. The spectra are normalized to unity. The simulations are performed for the |$\ket{\psi^{+}}$ polarisation state. Both data and MC samples are pre-filtered using the same selection conditions. The distributions are fitted with the function (red line) $f(\Delta\hat{\Phi})=a(1+b\cos(2\Delta\hat{\Phi}))$ with the reduced chi-2 equal to $\chi^{2}_{exp}=144.99/88\approx 1.65$ and $\chi^{2}_{MC}=114.93/88\approx 1.31$ for data and MC respectively. The MC distribution is in good agreement with the data distribution. The fitted values of the $a$ and $b$ parameters are equivalent within the statistical errors. Pull histogram shows the deviation between the fitted curve and histogram values expressed as $pull(x_{i})=\frac{H(x_{i})-f(x_{i})}{\sigma(x_{i})}$, where H is the histogram, the $f$ is the fitted function and $\sigma$ is an error associated with the given bin.
@@ Acknowledgments
The authors thank E. Czerwiński for his editorial remarks and suggestions. The authors thank P. Moskal and other J-PET collaboration members for providing the input data sample used in the analysis. BCH acknowledges gratefully that this research was funded in whole, or in part, by the Austrian Science Fund (FWF) project P36102-N. This work was completed with resources provided by the Świerk Computing Centre at the National Centre for Nuclear Research. For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
@@ Appendix B Calculation of the weighted mean of the visibility squared over the selection region
Let us consider a set of polarized two-photon pairs (e.g. in the initial state |$\ket{\Psi^{+}}$) measured by a perfect detector. For fixed scattering angles $(\tilde{\Theta}_{1},\tilde{\Theta}_{2})$, the conditional probability density function describing the measurement of the configuration $\Delta\hat{\Phi}$ will be given by:
$h(\Delta\hat{\Phi}|\tilde{\Theta}_{1},\tilde{\Theta}_{2})=\frac{1}{\pi}(1+\mathcal{V}(\tilde{\Theta}_{1})\mathcal{V}(\tilde{\Theta}_{2})\,\cos(2\Delta\hat{\Phi}))$ (18)
@@
which leads to:
$\langle h(\Delta\hat{\Phi})\rangle_{region}=\frac{1}{\pi}(1+\langle\mathcal{V}^{2}\rangle_{region}\cos(2\Delta\hat{\Phi}))$ (20)
The last equation can be interpreted as follows: if we form the $\Delta\hat{\Phi}$ distribution from all measured events in the selection area, and fit it with the function $f(\Delta\hat{\Phi})=a(1+b\cos(2\Delta\hat{\Phi}))$, then the $b$ parameter of the fit would correspond to the averaged visibility squared $\langle\mathcal{V}^{2}\rangle_{region}$ calculated over the selection region. Note that in this reasoning only $\Delta\hat{\Phi}$ is treated as a random variable, while $\tilde{\Theta}_{1},\tilde{\Theta}_{2}$ are treated as parameters, and the density $\rho(\tilde{\Theta}_{1},\tilde{\Theta}_{2})$ as fixed corresponding to the density of the measured events. The $\langle\mathcal{V}^{2}\rangle_{circle}$ factor is given by the equation:
[…]
where $C(\tilde{\Theta}_{a},\tilde{\Theta}_{b},R_{\tilde{\Theta}})$ is a circle defined by the inequality:
@@
To determine $\langle\mathcal{V}^{2}\rangle_{circle}$ from the data, one can use the discretized version of Eq. 21, and calculate the weighted mean over the selection region, where weighs are formed from the $n_{ij}$ corresponding to the registered number of events in the $i,j$-th bin: {dmath}V^2 ⟩_circle ≈∑_i=1^N_i∑_j=1^N_j δ(i,j,R_~Θ) nijNiNjV(~Θ_1(i))V(~Θ_2(j)) Δ~Θ_1Δ~Θ_2

### v2

@@ 5.4 Selection of high visibility region and detection efficiency corrections
[…]
@@ 5.5 Correlation plots
[…]
@@
[…]
@@ Acknowledgments
The authors thank E. Czerwiński for his editorial remarks and suggestions. The authors thank P. Moskal and other J-PET collaboration members for providing the input data sample used in the analysis. B.C.H. acknowledges gratefully that this research was funded in whole, or in part, by the Austrian Science Fund (FWF) project P36102-N (Grant DOI:10.55776/P36102). W.K. acknowledges the support by the Foundation for Polish Science through the FIRST TEAM FENG.02.02-IP.05-0152/23 programme co-financed by the European Union under the European Funds for Smart Economy 2021-2027 (FENG). This work was completed with resources provided by the Świerk Computing Centre at the National Centre for Nuclear Research. For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
@@ Appendix B Calculation of the weighted mean of the visibility squared over the selection region
[…]
$)h(\Delta\hat{\Phi}|\tilde{\Theta}_{1},\tilde{\Theta}_{2})=\frac{1}{\pi}(1+\mathcal{V}(\tilde{\Theta}_{1})\mathcal{V}(\tilde{\Theta}_{2})\,\cos(2\Delta\hat{\Phi}$)) (22)
@@
$\langle h(\Delta\hat{\Phi})\rangle_{region}=\int\limits_{region}\rho(\tilde{\Theta}_{1},\tilde{\Theta}_{2})h(\Delta\hat{\Phi}|\tilde{\Theta}_{1},\tilde{\Theta}_{2})d\tilde{\Theta}_{1}d\tilde{\Theta}_{2},$ (23)
which leads to:
$)\langle h(\Delta\hat{\Phi})\rangle_{region}=\frac{1}{\pi}(1+\langle\mathcal{V}^{2}\rangle_{region}\cos(2\Delta\hat{\Phi}$)) (24)
The last equation can be interpreted as follows: if we form the $\Delta\hat{\Phi}$ distribution from all measured events in the selection area, and fit it with the function $)f(\Delta\hat{\Phi})=a(1+b\cos(2\Delta\hat{\Phi}$)), then the $b$ parameter of the fit would correspond to the averaged visibility squared $\langle\mathcal{V}^{2}\rangle_{region}$ calculated over the selection region. Note that in this reasoning only $\Delta\hat{\Phi}$ is treated as a random variable, while $\tilde{\Theta}_{1},\tilde{\Theta}_{2}$ are treated as parameters, and the density $\rho(\tilde{\Theta}_{1},\tilde{\Theta}_{2})$ as fixed corresponding to the density of the measured events. The $\langle\mathcal{V}^{2}\rangle_{circle}$ factor is given by the equation:
$\langle\mathcal{V}^{2}\rangle_{circle}=\int\limits_{C(\tilde{\Theta}_{a},\tilde{\Theta}_{b},R_{\tilde{\Theta}})}\rho(\tilde{\Theta}_{1},\tilde{\Theta}_{2})\mathcal{V}(\tilde{\Theta}_{1})\mathcal{V}(\tilde{\Theta}_{2})d\tilde{\Theta}_{1}d\tilde{\Theta}_{2}$ (25)
@@
To determine $\langle\mathcal{V}^{2}\rangle_{circle}$ from the data, one can use the discretized version of Eq. 25, and calculate the weighted mean over the selection region, where weighs are formed from the $n_{ij}$ corresponding to the registered number of events in the $i,j$-th bin:
$\langle\mathcal{V}^{2}\rangle_{circle}\approx\sum_{i=1}^{N_{i}}\sum_{j=1}^{N_{j}}\delta(i,j,R_{\tilde{\Theta}})\frac{n_{ij}}{N_{i}N_{j}}\mathcal{V}(\tilde{\Theta}_{1}(i))\mathcal{V}(\tilde{\Theta}_{2}(j))\Delta\tilde{\Theta}_{1}\Delta\tilde{\Theta}_{2}$ (27)